# Prox-DBRO-VR: A Unified Analysis on Decentralized Byzantine-Resilient Composite Stochastic Optimization with Variance Reduction and Non-Asymptotic Convergence Rates

### v6

@@ V Experimental Results
Same-value attacks: As depicted in (4a), an $m=60$ multi-agent CPS consists of ${\left|\mathcal{R}\right|}=40$ reliable agents (yellow nodes) and $\left|\mathcal{B}\right|=20$ Byzantine agents (red nodes), where each Byzantine agent $o$, $o\in\mathcal{B}$, keeps sending ${z_{io,k}}=1000*{1_{n}}$ to its reliable neighbor $i$, $i\in{\mathcal{R}_{o}}$, at each iteration. Under this attack, the states of reliable agents can be easily blown up to sufficiently large values, which prevents the tested algorithms from convergence. Figs. (4b)-(4c) manifest that both Prox-DBRO-SAGA and Prox-DBRO-LSVRG achieve faster convergence than the other tested algorithms on the performance indices of OG and testing accuracy, while Prox-DBRO-LSVRG is slightly faster than Prox-DBRO-SAGA in this case. It can be found in Table IV that both Prox-DBRO-SAGA and Prox-DBRO-LSVRG attain also a smaller consensus error than the other tested algorithms at the final epoch. Note that the performance comparison takes no account of BRIDGE-B [23] due to its strict requirement of the number of its neighboring agents and high computational overhead. In a nutshell, the above simulation results demonstrate that Prox-DBRO-LSVRG and Prox-DBRO-SAGA achieve better convergence performance under different kinds of Byzantine attacks.
VI Conclusions
@@
In this paper, we proposed two decentralized Byzantine-resilient and variance-reduced stochastic gradient algorithms, namely Prox-DBRO-LSVRG and Prox-DBRO-SAGA, to resolve a category of non-smooth composite optimization problems over multi-agent CPSs in the presence of Byzantine agents. Theoretical analysis established both linear and sub-linear convergence rates for Prox-DBRO-LSVRG and Prox-DBRO-SAGA under different assumptions and parameter selections. In simulation, the proposed algorithms were applied to resolving a decentralized sparse soft-max regression task over multi-agent CPSs under different Byzantine attacks, which verifies the theoretical findings and demonstrates the better convergence performance of Prox-DBRO-LSVRG and Prox-DBRO-SAGA than the other notable decentralized algorithms. Future work will focus on extending the theoretical analysis of Prox-DBRO-LSVRG and Prox-DBRO-SAGA to the more general non-convex domain.
@@
According to Step 4 in Algorithm 3, at iteration $k$, $\forall k\geq 1$, the auxiliary variables $u_{i,k+1}^{l}$, $i\in\mathcal{R}$, take value $u_{i,k}^{l}$ or ${x_{i,k}}$, associated with probabilities $1-1/q_{i}$ and $1/q_{i}$, respectively. This observation is owing to the fact that selection of the random sample for Prox-DBRO-SAGA, at each iteration $k\geq 1$, is uniformly and independently executed. Hence, we have
@@
Recalling the definition of $t_{i,k}^{u}$ and combining Eqs. (15) and (17) give
@@
where the second inequality uses $1\leq{q_{\min}}\leq{q_{i}}\leq{q_{\max}}$, and the last equality is according to $f\left(x\right)=\sum\nolimits_{i\in\mathcal{R}}{{f_{i}}\left({{x_{i}}}\right)}$ and the definition of ${D_{F}}\left({{x_{k}},{x^{*}}}\right)$. Substituting the definition of $t_{k}^{u}$ obtains the relation (8). In view of Step 4 in Algorithm 2, we know that at iteration $k$, $\forall k\geq 1$, the auxiliary variables ${w_{i,k+1}}$, $i\in\mathcal{R}$, take value ${x_{i,k}}$ with probability $p_{i}$, or keep the most recent update ${w_{i,k}}$ with probability $1-{p_{i}}$. Therefore, it can be verified that
@@
Recalling the definition of $t_{i,k}^{w}$ and combining Eq. (20) with (21) give
${\mathbb{E}_{k}}\left[{t_{i,k+1}^{w}}\right]$ (22)
@@
$=$ $\left({1\!-{p_{i}}}\right){t_{i,k}^{w}}\!+\!{p_{i}}({{f_{i}}\left({{x_{i,k}}}\right)\!-\!{f_{i}}\left({{{\tilde{x}}^{*}}}\right)\!-\!\nabla{f_{i}}{{\left({{{\tilde{x}}^{*}}}\right)}^{\top}}\left({{x_{i,k}}\!-\!{{\tilde{x}}^{*}}}\right)}),$
@@
where the equality is due to the standard variance decomposition $\mathbb{E}_{k}\left[{\left\|A\right\|_{2}^{2}}\right]=\left\|{\mathbb{E}_{k}\left[A\right]}\right\|_{2}^{2}+\mathbb{E}_{k}\left[{\left\|{A-\mathbb{E}_{k}\left[A\right]}\right\|_{2}^{2}}\right]$, with $A={r_{i,k}^{u}}-\nabla{f_{i}}\left({{{\tilde{x}}^{*}}}\right)$. We continue to handle the first term in the right-hand-side of Eq. (23) as follows:
@@
where the second inequality utilizes $\mathbb{E}_{k}\left[{\left\|{B-\mathbb{E}_{k}\left[B\right]}\right\|_{2}^{2}}\right]\leq\mathbb{E}_{k}\left[{\left\|B\right\|_{2}^{2}}\right]$, with $B=\nabla f_{i}^{{s_{i,k}}}\left({u_{i,k}^{{s_{i,k}}}}\right)-\nabla f_{i}^{{s_{i,k}}}\left({{{\tilde{x}}^{*}}}\right)$, and the last equality applies the standard variance decomposition again. We proceed with substituting (24) into (23) to obtain
@@
${\mathbb{E}_{k}}\left[{\left\|{r_{k}^{u}-\nabla f\left({{{\tilde{x}}^{*}}}\right)}\right\|_{2}^{2}}\right]$ (26)
@@
Summing the both sides of (27) over $l$ from $1$ to $q_{i}$ becomes
@@
Since the local component function $f_{i}^{{s_{i,k}}}$, has a uniform distribution over the set $\left\{{f_{i}^{1},\ldots,f_{i}^{{q_{i}}}}\right\}$, it is natural to obtain
@@
Combining Eq. (29) and (28) and then summing over $i$ yield
@@
${\mathbb{E}_{k}}\left[{\left\|{r_{k}^{u}-\nabla f\left({{{\tilde{x}}^{*}}}\right)}\right\|_{2}^{2}}\right]$ (31)
@@
where we use the fact that ${f_{i}}\left({{x_{i,k}}}\right)=\left({1/{q_{i}}}\right)\sum\nolimits_{l=1}^{{q_{i}}}{f_{i}^{l}\left({{x_{i,k}}}\right)}$ and $f\left(x_{k}\right)=\sum\nolimits_{i\in\mathcal{R}}{{f_{i}}\left({{x_{i,k}}}\right)}$. Plugging (32) into (31) generates
${\mathbb{E}_{k}}\left[{\left\|{r_{k}^{u}-\nabla f\left({{{\tilde{x}}^{*}}}\right)}\right\|_{2}^{2}}\right]$ (33)
@@
Considering the $\mu$-strong convexity of the local objective function $f_{i}$, $\forall i\in\mathcal{R}$, we have
@@
Finally, one can obtain (10) via plugging the relation (34) into (33). For Prox-DBRO-LSVRG, we replace ${u_{i,k}^{l}}$ with ${{w_{i,k}}}$ to obtain (11), which completes the proof.
@@
${\mathbf{prox}}_{\alpha,G}\left\{x\right\}$ (35)
$=$ $\arg\mathop{\min}\limits_{y}\left\{{G\left(y\right)+\frac{1}{{2\alpha}}\left\|{y-x}\right\|_{2}^{2}}\right\}$
@@
which indicates ${\left[{\mathbf{prox}}_{\alpha,G}\left\{x\right\}\right]}_{i}={\mathbf{prox}}_{\alpha,g}\left\{x_{i}\right\}$. Based on this equality, it is straightforward to verify (12) with the help of the non-expansiveness of the proximal operator ${{\mathbf{prox}}_{\alpha,g}}$, which completes the proof.
@@
${0_{n}}\in{\nabla{f_{i}}\left({x_{i}^{*}}\right)+\partial_{x_{i}}g\left({x_{i}^{*}}\right)}+\frac{{{\phi_{i}}}}{2}\sum\limits_{j\in{\mathcal{R}_{i}}}{\partial{{\left\|{x_{i}^{*}-x_{j}^{*}}\right\|}_{a}}},\forall i\in\mathcal{R}.$ (36)
According to the definition of the sub-differential ${\partial}{\left\|{x_{i}^{*}-x_{j}^{*}}\right\|_{a}}=\left\{{{y_{ij}}\in{\mathbb{R}^{n}}|\left\langle{{y_{ij}},x_{i}^{*}}\right\rangle={{\left\|{x_{i}^{*}}\right\|}_{a}},{{\left\|{{y_{ij}}}\right\|}_{b}}\leq 1}\right\}$, there exist $g_{x_{i}}^{*}\in\partial_{x_{i}}g\left({x_{i}^{*}}\right)$ and ${\tilde{y}_{ij}}\in{\partial}{\left\|{x_{i}^{*}-x_{j}^{*}}\right\|_{a}}$, such that for $\forall i\in\mathcal{R}$
${\nabla{f_{i}}\left({x_{i}^{*}}\right)+g_{x_{i}}^{*}}+{\phi_{i}}\left({\sum\limits_{j\in{\mathcal{R}_{i}},i<j}{{{\tilde{y}}_{ij}}}-\sum\limits_{j\in{\mathcal{R}_{i}},i>j}{{{\tilde{y}}_{ji}}}}\right)=0_{n}.$ (37)
@@
${\nabla{f_{i}}\left({{{\tilde{x}}^{*}}}\right)+g_{\tilde{x}}^{*}}+{\phi_{i}}\left({\sum\limits_{j\in{\mathcal{R}_{i}},i<j}{{{\tilde{y}}_{ij}}}-\sum\limits_{j\in{\mathcal{R}_{i}},i>j}{{{\tilde{y}}_{ji}}}}\right)=0_{n},$ (38)
@@
where $g_{\tilde{x}}^{*}\in{\partial_{\tilde{x}}}g\left({{{\tilde{x}}^{*}}}\right)$. Since (38) can be decomposed into element-wise, without loss of generality, the rest proof assumes $n=1$, i.e., the scalar case. Via denoting $\Phi:={\text{diag}}\left\{\phi\right\}\in{\mathbb{R}^{\left|\mathcal{R}\right|\times\left|\mathcal{R}\right|}}$ with $\phi:={\left[{{\phi_{1}},{\phi_{2}},\ldots,{\phi_{\left|\mathcal{R}\right|}}}\right]^{\top}}\in{\mathbb{R}^{\left|\mathcal{R}\right|}}$, together with ${\psi_{i}}:={\nabla{f_{i}}\left({{{\tilde{x}}^{*}}}\right)+g_{\tilde{x}}^{*}}$, the task to prove (38) reduces to solving for a vector $\Psi$ with $\Psi:={\left[{\psi_{1},\psi_{2},\ldots,\psi_{\left|\mathcal{R}\right|}}\right]^{\top}}\in{\mathbb{R}^{\left|\mathcal{R}\right|}}$, such that the following relation holds
$\Phi\Pi{\tilde{y}}+\Psi=0_{\left|\mathcal{R}\right|},$ (39)
@@
where $\tilde{y}\in{\mathbb{R}^{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}$ is the collected form of $\tilde{y}_{ij}$ according to the order of edges in ${\mathcal{E}_{\mathcal{R}}}$. We need to solve for at least one solution $\tilde{y}$ meeting ${\left\|{{{\tilde{y}}_{ij}}}\right\|_{b}}\leq 1$ with $b>1$, such that (39) holds true. To proceed, we decompose the task into two parts.
@@
Part I: We first manifest that (39) has at least one solution. In view of the rank of the node-edge incidence matrix $\Pi$ is $\left|\mathcal{R}\right|-1$ and the null space of the columns is spanned by the all-one vector ${1_{\left|\mathcal{R}\right|}}$. Recalling the definition of ${\psi_{i}}$, the optimality condition of (1) is $\sum\nolimits_{i\in\mathcal{R}}{{\psi_{i}}}=0$. Therefore, we know that the columns of $\Pi$ and those of $\left[{\Phi\Pi,\Psi}\right]$ share the same null space, which indicates the same rank of $\Pi$ and $\left[{\Phi\Pi,\Psi}\right]$. The existence of solutions to (39) can be demonstrated according to the property of non-homogeneous linear equations.
@@
Part II: In this part, a solution with the $b$-norm of its elements no larger than $1$ is sought. Suppose that $y\in{\mathbb{R}^{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}$ is a solution to (39), such that $\Phi\Pi{y}+\Psi=0_{\left|\mathcal{R}\right|}$. We consider the least-squares solution $y=-{\Pi^{\dagger}}\Phi^{-1}\Psi$, where ${\Pi^{\dagger}}$ is the pseudo inverse of $\Pi$. Then, it suffices to prove that ${\left\|y\right\|_{b}}\leq 1$. Since ${\left\|y\right\|_{b}}={\left({\sum\nolimits_{i=1}^{\left|\mathcal{R}\right|}{{{\left|{{y_{i}}}\right|}^{b}}}}\right)^{1/b}},\forall b>1$, we know that ${\left\|y\right\|_{b}}\leq{\left\|y\right\|_{1}}$. Therefore, we derive
${\left\|y\right\|_{b}}\leq$ ${\left\|{{\Pi^{\dagger}}{\Phi^{-1}}\Psi}\right\|_{1}}$ (40)
$\leq$ ${\left\|{{\Pi^{\dagger}}}\right\|_{1}}{\left\|{{\Phi^{-1}}}\right\|_{1}}{\left\|\Psi\right\|_{1}}$
$\leq$ ${\left|\mathcal{R}\right|\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}\underline{\phi}^{-1}{\left\|{{\Pi^{\dagger}}}\right\|_{2}}{\left\|\Psi\right\|_{2}},$
@@
${\left\|y\right\|_{b}}\leq$ ${{{\lambda}_{\max}}}\left({{\Pi^{\dagger}}}\right)\frac{{{{\left|\mathcal{R}\right|\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}}}}{{{\underline{\phi}}}}{\left\|\Psi\right\|_{2}}$ (41)
$=$ $\frac{{{{\left|\mathcal{R}\right|\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}}}}{{{\underline{\phi}}{{{\lambda}_{\min}}\left(\Pi\right)}}}{\left\|\Psi\right\|_{2}}.$
@@
${\left\|y\right\|_{b}}\leq\frac{{{{\left|\mathcal{R}\right|}^{\frac{3}{2}}}\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}}{{{{{\lambda}_{\min}}\left(\Pi\right)}{\underline{\phi}}}}\mathop{\max}\limits_{i\in\mathcal{R}}\left|{{\psi_{i}}}\right|.$ (42)
@@
${\left\|y\right\|_{b}}\leq\frac{{{{\left|\mathcal{R}\right|}^{\frac{3}{2}}}\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}}{{{{{\lambda}_{\min}}\left(\Pi\right)}{\underline{\phi}}}}\mathop{\max}\limits_{i\in\mathcal{R}}{\left\|\nabla{f_{i}}\left({\tilde{x}^{*}}\right)+{g_{\tilde{x}}^{*}}\right\|_{\infty}}.$ (43)
The proof is completed by choosing an appropriate ${\underline{\phi}}$ to meet
$\frac{{{{\left|\mathcal{R}\right|}^{\frac{3}{2}}}\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}}{{{{{\lambda}_{\min}}\left(\Pi\right)}{\underline{\phi}}}}\mathop{\max}\limits_{i\in\mathcal{R}}{\left\|\nabla{f_{i}}\left({\tilde{x}^{*}}\right)+{g_{\tilde{x}}^{*}}\right\|_{\infty}}\leq 1.$
@@
Based on Proposition 2 and the optimality condition (36), we know that
@@
In view of the compact form (IV), it holds
${\mathbb{E}_{k}}\left[{\left\|{{x_{k+1}}-{x^{*}}}\right\|_{2}^{2}}\right]$ (45)

### v12

@@ V Numerical Experiments
Same-value attacks: As depicted in Fig. 4: LABEL:sub@Fig-3-1, an $m=60$ network consists of ${\left|\mathcal{R}\right|}=40$ reliable agents (yellow nodes) and $\left|\mathcal{B}\right|=20$ Byzantine agents (red nodes), where each Byzantine agent $b$, $b\in\mathcal{B}_{i}$, keeps sending ${z_{ib,k}}=1000*{1_{n}}$ to its reliable neighbor $i$, $i\in{\mathcal{R}}$, at each iteration. Under this attack, the states of reliable agents can be easily blown up to sufficiently large values, which prevents the tested algorithms from convergence. Figs. 4: LABEL:sub@Fig-3-2-LABEL:sub@Fig-3-3 manifest that the proposed algorithms achieve faster convergence than the other tested algorithms on the performance metrics of the optimal gap and testing accuracy, while Prox-DBRO-LSVRG is slightly faster than Prox-DBRO-SAGA in this case. It can be found in Table IV that the proposed algorithms attain also a smaller consensus error than the other tested algorithms at 150 epochs. Note that the consensus error of Prox-BRIDGE-B goes to a very large value since it adopts a vector-wise screening at each reliable agent, which results in a single surviving vector totally from one neighboring agent and thus easily leads to a large state variance between any two reliable agents. The performance comparison takes no account of BRIDGE-B [23] due to its strict requirement on the number of neighboring agents and high computational overhead. In a nutshell, the above numerical experiments demonstrate that while the proposed algorithms do not achieve the smallest consensus error only under zero-sum attacks, they achieve the best performance in all other cases.
(a) Optimal gap.
(b) Testing accuracy.
(c) Consensus error.
Figure 5: Performance degradation versus increasing proportions of Byzantine agents under sign-flipping attacks.
Sign-flipping attacks: In this case, we aim to verify the trade-off between the convergence accuracy and Byzantine resilience of the propose algorithms. The total number of agents including both reliable and Byzantine agents is $m=100$, where each Byzantine agent $b$, $b\in\mathcal{B}$, sends the falsified model ${z_{ib,k}}=-s_{b}\sum\nolimits_{j\in{\mathcal{R}_{i}}\cup\left\{i\right\}}{{x_{j,k}}}/\left({\left|{{\mathcal{R}_{i}}}\right|+1}\right)$ to their reliable neighbors $i$, $i\in{\mathcal{R}_{b}}$, where $s_{b}>0$ is the hyperparameter controlling the deviation of the attack. From Fig. 5, we can see that if the proportion or number of Byzantine agents increases, then the convergence accuracy regarding all three performance metrics becomes worse. This verifies the trade-off established in the theoretical result (see Theorems 2-3).
VI Conclusion
@@
In this paper, we proposed two Byzantine-resilient decentralized and VR stochastic gradient algorithms, namely Prox-DBRO-LSVRG and Prox-DBRO-SAGA, to resolve a category of nonsmooth composite finite-sum optimization problems over MASs in the presence of Byzantine agents. Theoretical analysis established both linear and sub-linear convergence rates for the proposed algorithms under different conditions. In the numerical experiments, the proposed algorithms were applied to resolving a decentralized sparse soft-max regression task over MASs under different Byzantine attacks, which verifies the theoretical findings and demonstrates the better convergence performance of the proposed algorithms than the other notable decentralized algorithms. However, both Prox-DBRO-LSVRG and Prox-DBRO-SAGA are not perfect, since they can only achieve exact sub-linear convergence in the absence of Byzantine agents. Future work will further investigate privacy issues and intermittent communication, which are also prevalent in MASs.
@@
According to Step 7 in Algorithm 3, at iteration $k$, $\forall k\geq 1$, the auxiliary variables $u_{i,k+1}^{l}$, $i\in\mathcal{R}$, take value $u_{i,k}^{l}$ or ${x_{i,k}}$, associated with probabilities $\left({1-1/{q_{i}}}\right)$ and $1/q_{i}$, respectively. This observation is owing to the fact that selection of the random sample for Prox-DBRO-SAGA, at each iteration $k\geq 1$, is uniformly and independently executed. Hence, we have
@@
Recalling the definition of $t_{i,k}^{u}$ and combining Eqs. (16) and (18), we obtain
@@
where the second inequality uses $1\leq{q_{\min}}\leq{q_{i}}\leq{q_{\max}}$, and the last equality is in accordance with the definition of ${D_{F}}\left({{x_{k}},{x^{*}}}\right)$. Substituting the definition of $t_{k}^{u}$ obtains the relation (8). In view of Step 7 in Algorithm 2, we know that at iteration $k$, $\forall k\geq 1$, the auxiliary variables ${w_{i,k+1}}$, $\forall i\in\mathcal{R}$, either take value ${x_{i,k}}$ with probability $p_{i}$, or keep the most recent update ${w_{i,k}}$ with probability $1-{p_{i}}$. Therefore, it can be verified that
@@
Recalling the definition of $t_{i,k}^{w}$ and combining Eq. (21) with (22), we have
${\mathbb{E}_{k}}\left[{t_{i,k+1}^{w}}\right]$
@@
$=$ $\left({1\!-{p_{i}}}\right){t_{i,k}^{w}}\!+\!{p_{i}}({{f_{i}}\left({{x_{i,k}}}\right)\!-\!{f_{i}}\left({{{\tilde{x}}^{*}}}\right)\!-\!\nabla{f_{i}}{{\left({{{\tilde{x}}^{*}}}\right)}^{\top}}\left({{x_{i,k}}\!-\!{{\tilde{x}}^{*}}}\right)}),$ (23)
@@
where the equality is due to the standard variance decomposition $\mathbb{E}_{k}\left[{\left\|A\right\|_{2}^{2}}\right]=\left\|{\mathbb{E}_{k}\left[A\right]}\right\|_{2}^{2}+\mathbb{E}_{k}\left[{\left\|{A-\mathbb{E}_{k}\left[A\right]}\right\|_{2}^{2}}\right]$, with $A={r_{i,k}^{u}}-\nabla{f_{i}}\left({{{\tilde{x}}^{*}}}\right)$. We continue to handle the first term in the RHS of Eq. (24) as follows:
@@
where the second inequality utilizes $\mathbb{E}_{k}\left[{\left\|{B-\mathbb{E}_{k}\left[B\right]}\right\|_{2}^{2}}\right]\leq\mathbb{E}_{k}\left[{\left\|B\right\|_{2}^{2}}\right]$, with $B=\nabla f_{i}^{{s_{i,k}}}\left({u_{i,k}^{{s_{i,k}}}}\right)-\nabla f_{i}^{{s_{i,k}}}\left({{{\tilde{x}}^{*}}}\right)$, and the last equality applies the standard variance decomposition again. We proceed to substitute (25) back into (24) to obtain
@@
${\mathbb{E}_{k}}\left[{\left\|{r_{k}^{u}-\nabla F\left({{{x}^{*}}}\right)}\right\|_{2}^{2}}\right]$ (27)
@@
Summing the both sides of (28) over $l$ from $1$ to $q_{i}$ obtains
@@
As the local component function $f_{i}^{{s_{i,k}}}$ has a uniform distribution over the set $\left\{{f_{i}^{1},\ldots,f_{i}^{{q_{i}}}}\right\}$, it is natural to obtain
@@
Combining Eq. (30) and (29), and then summing over $i$, yields
@@
${\mathbb{E}_{k}}\left[{\left\|{r_{k}^{u}-\nabla F\left({{{x}^{*}}}\right)}\right\|_{2}^{2}}\right]$ (32)
@@
where we use the fact that ${f_{i}}\left({{x_{i,k}}}\right)=\left({1/{q_{i}}}\right)\sum\nolimits_{l=1}^{{q_{i}}}{f_{i}^{l}\left({{x_{i,k}}}\right)}$ and $F\left(x_{k}\right)=\sum\nolimits_{i\in\mathcal{R}}{{f_{i}}\left({{x_{i,k}}}\right)}$. Plugging (33) back into (32) generates
${\mathbb{E}_{k}}\left[{\left\|{r_{k}^{u}-\nabla F\left({{{x}^{*}}}\right)}\right\|_{2}^{2}}\right]$ (34)
@@
Considering $\mu$-strong convexity of the local objective function $f_{i}$, $\forall i\in\mathcal{R}$, we have
@@
Finally, one can obtain (10) via plugging the relation (35) back into (34). For Prox-DBRO-LSVRG, we replace ${u_{i,k}^{l}}$ with ${{w_{i,k}}}$ to obtain (11), which completes the proof.
@@
${\mathbf{prox}}_{\alpha,G}\left\{x\right\}=$ $\arg\mathop{\min}\limits_{y}\left\{{G\left(y\right)+\frac{1}{{2\alpha}}\left\|{y-x}\right\|_{2}^{2}}\right\}$
@@
which indicates ${\left[{\mathbf{prox}}_{\alpha,G}\left\{x\right\}\right]}_{i}={\mathbf{prox}}_{\alpha,g}\left\{x_{i}\right\}$. Based on this equality, it is straightforward to verify (12) with the aid of the non-expansiveness of the proximal operator ${{\mathbf{prox}}_{\alpha,g}}$, which completes the proof.
@@
${0_{n}}\in{\nabla{f_{i}}\left({x_{i}^{*}}\right)+\partial g\left({x_{i}^{*}}\right)}+\frac{{{\phi}}}{2}\sum\limits_{j\in{\mathcal{R}_{i}}}{\partial{{\left\|{x_{i}^{*}-x_{j}^{*}}\right\|}_{a}}},\forall i\in\mathcal{R}.$ (37)
According to the definition of the subdifferential ${\partial}{\left\|{x_{i}^{*}-x_{j}^{*}}\right\|_{a}}=\left\{{{y_{ij}}\in{\mathbb{R}^{n}}|\left\langle{{y_{ij}},x_{i}^{*}}\right\rangle={{\left\|{x_{i}^{*}}\right\|}_{a}},{{\left\|{{y_{ij}}}\right\|}_{b}}\leq 1}\right\}$, there exist $g^{\prime}\left({x_{i}^{*}}\right)\in\partial g\left({x_{i}^{*}}\right)$ and ${\tilde{y}_{ij}}\in{\partial}{\left\|{x_{i}^{*}-x_{j}^{*}}\right\|_{a}}$, such that for $i\in\mathcal{R}$
${\nabla{f_{i}}\left({x_{i}^{*}}\right)+g^{\prime}\left({x_{i}^{*}}\right)}+{\phi}\left({\sum\limits_{j\in{\mathcal{R}_{i}},i<j}{{{\tilde{y}}_{ij}}}-\sum\limits_{j\in{\mathcal{R}_{i}},i>j}{{{\tilde{y}}_{ji}}}}\right)=0_{n}.$ (38)
@@
${\nabla{f_{i}}\left({{{\tilde{x}}^{*}}}\right)+g^{\prime}\left({{{\tilde{x}}^{*}}}\right)}+{\phi}\left({\sum\limits_{j\in{\mathcal{R}_{i}},i<j}{{{\tilde{y}}_{ij}}}-\sum\limits_{j\in{\mathcal{R}_{i}},i>j}{{{\tilde{y}}_{ji}}}}\right)=0_{n},$ (39)
@@
where $g^{\prime}\left({{{\tilde{x}}^{*}}}\right)\in{\partial}g\left({{{\tilde{x}}^{*}}}\right)$. Since (39) can be decomposed into element-wise, without loss of generality, the rest proof assumes $n=1$, i.e., the scalar case. Via denoting $\Psi:={\mathbf{col}}{\left\{{{\psi_{i}}}\right\}_{i\in\mathcal{R}}}\in{\mathbb{R}^{\left|\mathcal{R}\right|}}$ with ${\psi_{i}}:={\nabla{f_{i}}\left({{{\tilde{x}}^{*}}}\right)+g^{\prime}\left({{{\tilde{x}}^{*}}}\right)}$, the task to prove (39) reduces to solving for a vector ${\tilde{y}}$, such that the following relation holds
$\phi\Pi{\tilde{y}}+\Psi=0_{\left|\mathcal{R}\right|},$ (40a)
${\left\|{{{\tilde{y}}}}\right\|_{b}}\leq 1,$ (40b)
@@
where $\tilde{y}\in{\mathbb{R}^{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}$ is the collected form of $\tilde{y}_{ij}$ according to the order of edges in ${\mathcal{E}_{\mathcal{R}}}$. We need to solve for at least one solution $\tilde{y}$ meeting ${\left\|{{{\tilde{y}}}}\right\|_{b}}\leq 1$ with $b>1$, such that (40a) holds true. To proceed, we decompose the task into two parts.
@@
Part I: We first manifest that (40a) has at least one solution. In view of the rank of the node-edge incidence matrix $\Pi$ is $\left|\mathcal{R}\right|-1$ and the null space of the columns is spanned by the all-one vector ${1_{\left|\mathcal{R}\right|}}$. Recalling the definition of ${\psi_{i}}$, the optimality condition of (1) is $\sum\nolimits_{i\in\mathcal{R}}{{\psi_{i}}}=0$. Therefore, we know that the columns of $\Pi$ and those of $\left[{\phi\Pi,\Psi}\right]$ share the same null space, which indicates the same rank of $\Pi$ and $\left[{\phi\Pi,\Psi}\right]$ according to the rank-nullity theorem. The existence of solutions to (40) can be demonstrated according to the property of non-homogeneous linear equations.
@@
Part II: We provide a specific condition to ensure that the solution to (40a) satisfies (40b) as well. Suppose that $y\in{\mathbb{R}^{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}$ is a solution to (40a), such that $\phi\Pi{y}+\Psi=0_{\left|\mathcal{R}\right|}$. We consider the least-squares solution $y=-{\Pi^{\dagger}}\Psi/\phi$, where ${\Pi^{\dagger}}$ is the Moore-Penrose pseudo-inverse of $\Pi$. Then, it suffices to prove that ${\left\|y\right\|_{b}}\leq 1$. Since ${\left\|y\right\|_{b}}={\left({\sum\nolimits_{i=1}^{\left|\mathcal{R}\right|}{{{\left|{{y_{i}}}\right|}^{b}}}}\right)^{1/b}},\forall b>1$, we know that ${\left\|y\right\|_{b}}\leq{\left\|y\right\|_{1}}$. Therefore, we derive
${\left\|y\right\|_{b}}\leq$ $\frac{1}{\phi}{\left\|{{\Pi^{\dagger}}\Psi}\right\|_{1}}$ (41)
$\leq$ $\frac{1}{\phi}{\left\|{{\Pi^{\dagger}}}\right\|_{1}}{\left\|\Psi\right\|_{1}}$
$\leq$ $\frac{1}{\phi}{\left|\mathcal{R}\right|\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}{\left\|{{\Pi^{\dagger}}}\right\|_{2}}{\left\|\Psi\right\|_{2}},$
@@
${\left\|y\right\|_{b}}\leq$ ${{{\lambda}_{\max}}}\left({{\Pi^{\dagger}}}\right)\frac{{{{\left|\mathcal{R}\right|\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}}}}{{{\phi}}}{\left\|\Psi\right\|_{2}}$ (42)
$=$ $\frac{{{{\left|\mathcal{R}\right|\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}}}}{{{\phi}{{{\lambda}_{\min}}\left(\Pi\right)}}}{\left\|\Psi\right\|_{2}}.$
@@
${\left\|y\right\|_{b}}\leq\frac{{{{\left|\mathcal{R}\right|}^{\frac{3}{2}}}\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}}{{{{{\lambda}_{\min}}\left(\Pi\right)}{\phi}}}\mathop{\max}\limits_{i\in\mathcal{R}}\left|{{\psi_{i}}}\right|.$ (43)
@@
${\left\|y\right\|_{b}}\leq\frac{{{{\left|\mathcal{R}\right|}^{\frac{3}{2}}}\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}}{{{{{\lambda}_{\min}}\left(\Pi\right)}{\phi}}}\mathop{\max}\limits_{i\in\mathcal{R}}{\left\|\nabla{f_{i}}\left({\tilde{x}^{*}}\right)+{g^{\prime}\left({{{\tilde{x}}^{*}}}\right)}\right\|_{\infty}}.$ (44)
The proof is completed by choosing an appropriate ${\phi}$ to meet
$\frac{{{{\left|\mathcal{R}\right|}^{\frac{3}{2}}}\sqrt{\left|{{\mathcal{E}_{\mathcal{R}}}}\right|}}}{{{{{\lambda}_{\min}}\left(\Pi\right)}{\phi}}}\mathop{\max}\limits_{i\in\mathcal{R}}{\left\|\nabla{f_{i}}\left({\tilde{x}^{*}}\right)+{g^{\prime}\left({{{\tilde{x}}^{*}}}\right)}\right\|_{\infty}}\leq 1,$
which provides the condition to ensure the validity of (40) and thus finishes the proof.
@@
Without loss of generality, we let ${{\tilde{y}}_{ji}}=-{{\tilde{y}}_{ij}}$ for $j>i$ such that by defining ${{\tilde{y}}_{i}}:=\sum\nolimits_{j\in{{\mathcal{R}}_{i}}}{{{\tilde{y}}_{ij}}}$ and ${\partial_{x}}\chi\left({{x^{*}}}\right):={\mathbf{col}}{\left\{{{{\tilde{y}}_{i}}}\right\}_{i\in\mathcal{R}}}\in{\mathbb{R}^{\left|\mathcal{R}\right|n}}$, it follows from the optimality condition (37) that
@@
In view of the compact form (7), it holds
${\mathbb{E}_{k}}\left[{\left\|{{x_{k+1}}-{x^{*}}}\right\|_{2}^{2}}\right]$